A ball is batted into the air and caught at a point 100 m distance horizontally in 4.0 seconds. Neglecting air resistance. Find the maximum height attain by the ball.
1
Expert's answer
2020-09-24T11:09:35-0400

The height of a body thrown at some angle above horizontal:


"h=\frac{v^2\text{sin}^2\theta}{2g}."

The range is


"R=\frac{v^2\text{ sin}(2\theta)}{g},\\\space\\
\text{ sin}(2\theta)=2\text{cos}\theta\text{ sin}\theta,\\\space\\
R=\frac{2v^2\text{cos}\theta\text{ sin}\theta}{g}."

On the other hand, the range is


"R=v\text{ cos}\theta \cdot t,\\\space\\
v\text{ cos}\theta =\frac{R}{t}."

Substitute to the equation with sine of a double angle:

"R=\frac{2Rv\text{ sin}\theta}{gt},\\\space\\
v\text{ sin}\theta=\frac{gt}{2}."

Substitute this into the first equation for height:


"h=\frac{(gt/2)^2}{2g}=\frac{gt^2}{8}=19.6\text{ m}."